# Good involutions of conjugation subquandles

### v4

@@ 8.3. Enumeration
We implemented Algorithms 8.1 and 8.5 in GAP [GAP4] and used them to compute $\operatorname{Good}(\operatorname{Conj}G)$ and $\operatorname{Good}(\operatorname{Core}G)$ for various groups $G$ up to order $23$. We provide our code and raw data (including the explicit mappings $\rho:G\to G$) in a GitHub repository [code].
Table 2 counts good involutions of conjugation quandles $\operatorname{Conj}G$ of nonabelian groups up to order $22$. Table 3 does the same for core quandles $\operatorname{Core}G$ of groups of exponent greater than $2$ up to order $23$. In Appendix B, we specify these counts for each group $G$ considered in Tables 2 and 3.
Table 2. Number of good involutions across all nontrivial conjugation quandles $\operatorname{Conj}G$ of groups $G$ up to order $22$. Cf. Table 6.
@@
Table 3. Number of good involutions across all nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order $23$. Cf. Table LABEL:tab:app2.
@@ 9. Constructions
Carrying over our notation from Section 6, let $G$ be a group, let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$, and let $H:=\langle X\rangle\leq G$. Also, let $\iota:G\to G$ be the inversion map $g\mapsto g^{-1}$.
@@ 9.1. Constructing CNS-quandles
  • There exists an element $x\in X$ such that $x^{2}\notin Z(H)$.
@@
By Lemma 2.10, $\operatorname{Conj}X$ is an inversion-closed subquandle of $\operatorname{Conj}G$, so Example 3.12 states that $\iota|_{X}$ is a good involution of $\operatorname{Conj}X$. Also, Lemma 2.25 states that $\operatorname{Conj}X$ is connected, and Observation 2.11 states that $\operatorname{Conj}X$ is noninvolutory. This completes the proof. ∎
@@ 9.2. Examples
Let $G$ be the special linear group $\mathrm{SL}(2,5)$. We study good involutions of several connected subquandles of $\operatorname{Conj}G$. Using GAP [GAP4], we find that $G$ has seven conjugacy classes $X$ of size greater than $1$: four of size $12$, two of size $20$, and one of size $30$.
@@
It follows that each of the seven aforementioned conjugacy classes $X$ generates $G$, so by Lemma 2.25, all of them are connected subquandles $\operatorname{Conj}X$ of $\operatorname{Conj}G$. Using GAP, we verify that the subquandles of orders $12$ and $20$ satisfy the conditions in Proposition 9.1. Therefore, $\operatorname{Conj}G$ contains four CNS-subquandles $(\operatorname{Conj}X,\iota|_{X})$ of order $12$ and two CNS-subquandles of order $20$.
On the other hand, the subquandle $\operatorname{Conj}X$ of order $30$ satisfies only the first two conditions in Proposition 9.1, so $(\operatorname{Conj}X,\iota|_{X})$ is a connected symmetric kei. Since the element $-I_{2}\in Z(G)$ corresponds to the identity map in Corollary 6.11, it follows from Proposition 3.15 and Corollary 6.11 that
@@
because $p>2$. Hence, Proposition 9.1 states that $(\operatorname{Conj}X,\iota|_{X})$ is a CNS-quandle. In particular, the formula for $|X|$ given in [fulman]*Sec. 2 yields the first part of Proposition 1.8.
Finally, we show that $\iota|_{X}$ is the only good involution of $\operatorname{Conj}X$. By Observation 6.10 and Corollary 6.11, it suffices to show that there exists a transvection $B\in X$ such that for all transvections $A\in X$, the product $AB$ does not equal any nonidentity element $\lambda I$ of $Z(H)$. But this follows from the fact that $A$ and $B$ are unipotent (due to being transvections).
@@ 9.2.1. An infinite-order example
Let $G$ be the special unitary group $\mathrm{SU}(2)$. Motivated by the knot-theoretic and representation-theoretic applications of connected subquandles of $\operatorname{Conj}G$citelist[su2][su2-2], we study good involutions of these subquandles.
@@
with $\theta\in[0,\pi]$ represent all conjugacy classes of $G$, and an element $g\in G$ is conjugate to $d_{\theta}$ if and only if the eigenvalues of $g$ are $e^{\pm i\theta}$. Let $X_{\theta}$ be the conjugacy class of $d_{\theta}$ in $G$.
Recall that the only nontrivial normal subgroup of $G$ is the center
$Z(G)=\{\pm I_{2}\}.$
Fix $\theta\in(0,\pi)$. It follows that $X_{\theta}$ generates $G$, so by Lemma 2.25, $\operatorname{Conj}X_{\theta}$ is connected. By the above discussion, $\operatorname{Conj}X_{\theta}$ is also inversion-closed. By verifying the conditions in Corollary 6.11 and Proposition 9.1, we deduce that
@@ 10.1.1. Corollaries
Let $n\in\mathbb{Z}^{+}$ be a positive integer, and define affine transformations $\alpha_{m,b}:\mathbb{Z}/2n\mathbb{Z}\to\mathbb{Z}/2n\mathbb{Z}$ by $x\mapsto mx+b$. Let $R_{2n}=T(\mathbb{Z}/2n\mathbb{Z})$ be the dihedral quandle of order $2n$. Then
@@ 10.1.2. Computer search
Table 5. Enumeration of isomorphism classes of various types of Legendrian racks up to order $8$.
@@
Adding to our classification of symmetric racks up to order $8$ using GAP [GAP4], we used the classification of GL-racks up to order $8$ in [taGL] to classify various types of Legendrian racks up to isomorphism.
@@ 10.2.3. Medial racks
We discuss a class of racks called medial racks or abelian racks, which Joyce [joyce] introduced to study Alexander invariants of knots. [^4] In this paper, we adopt the name “medial” over “abelian.” This is to prevent confusion with commutative racks, which satisfy the much rarer condition that $s_{x}(y)=s_{y}(x)$ for all $x,y\in X$. Certain categorical and universal-algebraic notions (see, for example, [rack-roll]*Sec. 3, [Hom]*Thm. 12, or [taGL]*Sec. 2.3.3) allow medial racks to enhance coloring invariants of smooth knots [Hom]*Ex. 9 and Legendrian knots [taGL]*Prop. 7.2.
@@ 10.3. Preliminaries: GL-racks
We discuss GL-racks, which Karmakar et al. [karmakarGL] and Kimura [bi] independently introduced in 2023 to develop invariants of Legendrian knots in contact three-space. See [taGL] for a development of the algebraic theory of GL-racks.
@@ 10.3.1. Centrality of GL-structures
By Definition 10.16, all integer powers of GL-structures $\mathtt{u}$ and canonical rack automorphisms $\theta$ lie in the categorical center $Z(\mathsf{GLR})$. In fact, $Z(\mathsf{GLR})\cong\mathbb{Z}^{2}$ is the free abelian group generated by these two collections of automorphisms; see [taGL]*Thm. 4.16.
@@ Appendix B Census of good involutions
In this section, we enumerate good involutions of nontrivial conjugation quandles $\operatorname{Conj}G$ and nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order $23$.
@@
  • Let $D_{n}$ and $\operatorname{Dic}_{n}$ denote the dihedral group of order $2n$ and the dicyclic group of order $4n$, respectively.
  • Let $Q_{8}$, $QD_{16}$, and $M_{16}$ denote the quaternion group of order $8$, the quasidihedral group of order $16$, and the modular maximal-cyclic group of order $16$, respectively.
In Table 6, for all nonabelian groups $G$ up to order $22$, we compare the number of good involutions of $\operatorname{Conj}G$ against the upper bound on this number from Corollary 6.12. These results motivate Problems 11.6 and 11.7. We only consider nonabelian groups because conjugation quandles of abelian groups are trivial. Cf. Table 2.
Table 6. Number of good involutions of nontrivial conjugation quandles $\operatorname{Conj}G$ of groups $G$ up to order $22$, compared against the upper bound from Corollary 6.12.
Group Order $|{\operatorname{Good}(\operatorname{Conj}G)}|$ $|Z(G)|^{k(G)}$
@@
In Table LABEL:tab:app2, for all groups $G$ of exponent greater than $2$ up to order $23$, we compare the number of good involutions of $\operatorname{Core}G$ against the upper bound on this number from Corollary 7.7. We do not consider elementary abelian $2$-groups because their core quandles are trivial. Cf. Table 3.
@@
Table 7. Number of good involutions of nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order $23$, compared against the upper bound from Corollary 7.7.
Group Order $|{\operatorname{Good}(\operatorname{Core}G)}|$ $|A|^{|\mathcal{O}|}$
@@
$\mathbb{Z}/6\times\mathbb{Z}/3\mathbb{Z}$ 18 2 4

### v5

@@ 8.3. Enumeration
We implemented Algorithms 8.1 and 8.5 in GAP [GAP4] and used them to compute $\operatorname{Good}(\operatorname{Conj}G)$ and $\operatorname{Good}(\operatorname{Core}G)$ for various groups $G$ up to order 232323. We provide our code and raw data (including the explicit mappings $\rho:G\to G$) in a GitHub repository [code].
Table 2 counts good involutions of conjugation quandles $\operatorname{Conj}G$ of nonabelian groups up to order 222222. Table 3 does the same for core quandles $\operatorname{Core}G$ of groups of exponent greater than 222 up to order 232323. In Appendix B, we specify these counts for each group $G$ considered in Tables 2 and 3.
Table 2. Number of good involutions across all nontrivial conjugation quandles $\operatorname{Conj}G$ of groups $G$ up to order 222222. Cf. Table 6.
@@
Table 3. Number of good involutions across all nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order 232323. Cf. Table LABEL:tab:app2.
@@ 9. Constructions
Carrying over our notation from Section 6, let $G$ be a group, let $\operatorname{Conj}X$ be a subquandle of $\operatorname{Conj}G$, and let $H:=\langle X\rangle\leq G$. Also, let $\iota:G\to G$ denote the inversion map $g\mapsto g^{-1}$.
@@ 9.1. Constructing CNS-quandles
  • There exists an element $x\in X$ such that $x^{2}\notin Z(H)$.
@@
This follows from Lemma 2.25, Observation 2.11, and Example 3.13. ∎
@@ 9.2. Examples
Let $G$ be the special linear group $\mathrm{SL}(2,5)$. We study good involutions of several connected subquandles of $\operatorname{Conj}G$. Using GAP [GAP4], we find that $G$ has seven conjugacy classes $X$ of size greater than 111: four of size 121212, two of size 202020, and one of size 303030.
@@
It follows that each of the seven aforementioned conjugacy classes $X$ generates $G$, so by Lemma 2.25, all of them are connected subquandles $\operatorname{Conj}X$ of $\operatorname{Conj}G$. Using GAP, we verify that the subquandles of orders 121212 and 202020 satisfy the conditions in Proposition 9.1. Therefore, $\operatorname{Conj}G$ contains four CNS-subquandles $(\operatorname{Conj}X,\iota|_{X})$ of order 121212 and two CNS-subquandles of order 202020.
On the other hand, the subquandle $\operatorname{Conj}X$ of order 303030 satisfies only the first two conditions in Proposition 9.1, so $(\operatorname{Conj}X,\iota|_{X})$ is a connected symmetric kei. Since the element $-I_{2}\in Z(G)$ corresponds to the identity map in Corollary 6.11, it follows from Proposition 3.16 and Corollary 6.11 that
@@
[…]
Finally, we show that $\iota|_{X}$ is the only good involution of $\operatorname{Conj}X$. By Corollary 6.11 and Example 6.8, it suffices to show that there exists a transvection $B\in X$ such that for all transvections $A\in X$, the product $AB$ does not equal any nonidentity element $\lambda I$ of $Z(H)$. But this follows from the fact that $A$ and $B$ are unipotent (due to being transvections).
@@ 9.2.1. An infinite-order example
[…]
@@
with $\theta\in[0,\pi]$ represent all conjugacy classes of $G$, and an element $g\in G$ is conjugate to $d_{\theta}$ if and only if the eigenvalues of $g$ are $e^{\pm i\theta}$. Let $X_{\theta}$ denote the conjugacy class of $d_{\theta}$ in $G$.
Fix $\theta\in(0,\pi)$. Since the only nontrivial normal subgroup of $G$ is the center
$Z(G)=\{\pm I_{2}\},$
it follows that $X_{\theta}$ generates $G$. Thus, Lemma 2.25 shows that $\operatorname{Conj}X_{\theta}$ is connected. By the above discussion, $\operatorname{Conj}X_{\theta}$ is also inversion-closed. By verifying the conditions in Corollary 6.11 and Proposition 9.1, we deduce that
@@ 10.1.1. Corollaries
[…]
@@ 10.1.2. Computer search
Table 5. Enumeration of isomorphism classes of various types of Legendrian racks up to order 888.
@@
Adding to our classification of symmetric racks up to order 888 using GAP [GAP4], we used the classification of GL-racks up to order 888 in [taGL] to classify various types of Legendrian racks up to isomorphism.
@@ 10.2.3. Medial racks
We discuss a class of racks called medial racks or abelian racks, which Joyce [joyce] introduced to study Alexander invariants of knots. [^4] In this paper, we adopt the name “medial” over “abelian.” This is to prevent confusion with commutative racks, which satisfy the much rarer condition that $s_{x}(y)=s_{y}(x)$ for all $x,y\in X$. Certain categorical and universal-algebraic notions (see, for example, [rack-roll]*Sec. 3, [Hom]*Thm. 12, or [taGL]*Sec. 2.3.3) allow medial racks to enhance coloring invariants of smooth knots [Hom]*Ex. 9 and Legendrian knots [taGL]*Cor. 3.21.
@@ 10.3. Preliminaries: GL-racks
We discuss GL-racks, which Karmakar et al. [karmakarGL] and Kimura [bi] independently introduced in 2023 to develop invariants of Legendrian knots in contact three-space. See [taGL] for a development of the algebraic theory of GL-racks and [he] for applications.
@@ 10.3.1. Centrality of GL-structures
By Definition 10.16, all integer powers of GL-structures $\mathtt{u}$ and canonical rack automorphisms $\theta$ lie in the categorical center $Z(\mathsf{GLR})$. In fact, $Z(\mathsf{GLR})\cong\mathbb{Z}^{2}$ is the free abelian group generated by these two collections of automorphisms; see [taGL]*Thm. 5.1.
@@ Appendix B Census of good involutions
In this section, we enumerate good involutions of nontrivial conjugation quandles $\operatorname{Conj}G$ and nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order 232323.
@@
[…]
  • Let $Q_{8}$, $QD_{16}$, and $M_{16}$ denote the quaternion group of order 888, the quasidihedral group of order 161616, and the modular maximal-cyclic group of order 161616, respectively.
In Table 6, for all nonabelian groups $G$ up to order 222222, we compare the number of good involutions of $\operatorname{Conj}G$ against the upper bound on this number from Corollary 6.12. These results motivate Problems 11.6 and 11.7. We only consider nonabelian groups because conjugation quandles of abelian groups are trivial. Cf. Table 2.
Table 6. Number of good involutions of nontrivial conjugation quandles $\operatorname{Conj}G$ of groups $G$ up to order 222222, compared against the upper bound from Corollary 6.12.
Group Order $||{\operatorname{Good}(\operatorname{Conj}G)}|$$|Z(G)|^{k(G)}$
@@
In Table LABEL:tab:app2, for all groups $G$ of exponent greater than 222 up to order 232323, we compare the number of good involutions of $\operatorname{Core}G$ against the upper bound on this number from Corollary 7.14. We do not consider elementary abelian 222-groups because their core quandles are trivial. Cf. Table 3.
@@
Table 7. Number of good involutions of nontrivial core quandles $\operatorname{Core}G$ of groups $G$ up to order 232323, compared against the upper bound from Corollary 7.14.
Group Order $||{\operatorname{Good}(\operatorname{Core}G)}|$$|T|^{|\mathcal{O}|}$
@@
$\mathbb{Z}/6\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$ 18 2 4